OpenStudy (anonymous):

Which equation would best help solve the following problem? Petra throws a ball up with an initial vertical velocity of 35 ft/s from a platform that is at ground level. How long will the ball be in the air? A. 16t2 + 35t = 0 B. –4.9t2 – 35t = 0 C. –16t2 – 35t = 0 D. –16t2+ 35t = 0

OpenStudy (anonymous):

@pinkbubbles

OpenStudy (anonymous):

deeezzz nuts loool

OpenStudy (anonymous):

loool

OpenStudy (anonymous):

h=−16t2+vi+h0 vi initial velocity h0 initial height

OpenStudy (anonymous):

D

OpenStudy (anonymous):

16t2 + 35t = 0

OpenStudy (anonymous):

so your saying its A?

OpenStudy (anonymous):

@pinkbubbles

Vocaloid (vocaloid):

take another look at Deez's answer

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

Deeezz nutsss

OpenStudy (anonymous):

Got em2

OpenStudy (anonymous):

YEEES

OpenStudy (welshfella):

use one of the equations for constant acceleration s = 0 , u = 35 , a = -32, t = time s = ut + 0.5at^2 0 = 35t - 0.5*32t^2 35t - 16t^2 = 0

OpenStudy (welshfella):

D

OpenStudy (welshfella):

s = displacement , u = initial velocity, a = acceleration due to gravity , t = time